Express the confidence interval 0.466<p<0.5560.466<p<0.556 in the form of ˆp±MEp^±ME.

±±
Answer #1

Here the given confidence interval is 0.466 < p < 0.556

Here width of the interval = 0.556 - 0.466 = 0.09 ........(1)

we know that the width of the interval = 2 (margin of error) = 2 (M.E) .........(2)

here from (1) & (2)

2 (M.E) = 0.09

M.E = 0.09/2

M.E = 0.045

now = Upper limit - M.E = 0.556 - 0.045 = 0.511

The confidence interval in the form of is